The question is as follows:
The Fed purchases $12 million in US bonds from a bond dealer and the dealer's bank credits the dealer's account. The required reserve ratio is 13 percent. ......calculate how much will the bank be able to lend to its customers following the Fed's purchase.
I have total Reserves - required Reserves = excess reserves = loans to its customers.......
or
12 - 1.32 = 10.68 billion.

This came from the "show me how to solve the problem"........But, where does the 1.32 come from?

Can you please help me?

Thank you!!

Sara Wright
1
Expert's answer
2013-12-26T13:18:47-0500
This calculation is incorrect. The right answer should be: 12 - 1.56 = 10.44, because the required reserves = 12*0.13 = 1.56